# Every finite set of natural numbers is realizable as algebraic periods of a Morse–Smale diffeomorphism

### v1

@@ 3.1. Existence of periodic points for Morse–Smale diffeormophisms and transversal maps
To formulate a consequences of our main result related to the existence of periodic points we remind the notion of a transversal map.
Definition 3.5 ([Do], [Fr]).
Let $f\colon{\mathcal{U}}\to M$ be a $C^{1}$-map of an open subset ${\mathcal{U}}$ of a manifold $M$, $\dim M=d$. We say that $f$ is transversal if for any $m\in{\mathbb{N}}$ and $x\in P^{m}(f)$, $1\notin\sigma(Df^{m}(x))$, where $\sigma$ denotes the spectrum of the derivative $D$ of $f^{m}(x)$.
The set of all transversal maps ${\mathcal{U}}\to M$ is denoted by $C_{T}({\mathcal{U}},M)$ or $C_{T}(M)$ if ${\mathcal{U}}=M$. The main property of the class of transversal maps is given in the following theorem (see [Do] if $M=\mathbb{R}^{d}$, and [Seidel] for the general case, also [Je-Ma] for an exposition).
@@
The set $C^{\infty}_{T}({\mathcal{U}},M)$ is generic in $C^{0}({{\mathcal{U}}},M)$, i.e. it is the intersection $C^{\infty}_{T}({\mathcal{U}},M)=\bigcap_{n=1}^{\infty}G_{n}$ where $G_{n}$ is open and dense in $C^{0}({{\mathcal{U}}},M)$ In particular every map $f\colon{\mathcal{U}}\to M$ is homotopic to a transversal map $h\colon{\mathcal{U}}\to M$.
A geometric property of a transversal map states the following.
@@
Let $M$ be a closed manifold. For any $f\in C_{T}(M)$ and every $m\in\mathbb{N}$ the set $P^{m}(f)$ consists of isolated points.
By definition, Morse–Smale diffeomorphisms belong to $C_{T}(M)$ (cf. [Smale], [Fr]). We denote the set of transversal self-diffeomorphisms of $M$ by $DC_{T}(M)$ and the set of Morse–Smale diffeomorphisms by $DC_{MS}(M)$.
Finally, all mappings in $C_{T}(M)$, thus $DC_{T}(M)$ and $DC_{MS}(M)$, have the following geometric property, which states that non-vanishing $a_{n}(f)$, $n$-odd, implies that $P_{n}(f)\neq\emptyset$, i.e. $n\in Per(f)$. We have (cf. [Do], and [Fr] for Morse–Smale diffeomorphism).
Proposition 3.8 (see [Je-Ma, Corollary 3.3.10] for an exposition).
@@
Let $f:M\to M$ be a transversal map. If $a_{n}(f)\neq 0\,$ then $\;\begin{cases}P_{n}(f)\cup P_{\frac{n}{2}}(f)\neq\emptyset\;\,{\text{if}}\;n\;{\text{is even}},\cr P_{n}(f)\neq\emptyset\;\;{\text{if}}\;n\;{\text{is odd}}.\end{cases}$
The above proposition allows us to formulate the analytical (dynamical) consequence of Theorem 1.4. Put ${\mathcal{A}P}_{odd}(f)={\mathcal{A}P}(f)\cap\{2\mathbb{N}-1\}$, and ${\mathcal{A}P}_{ev}(f)={\mathcal{A}P}(f)\cap\{2\mathbb{N}\}$.
@@
Let ${\mathcal{A}}\subset\mathbb{N}$ be a finite subset of natural numbers and $f\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ be a preserving, or reversing orientation Morse–Smale diffeomorphism of orientable surface of genus $\mathrm{g}$ given by Theorem 1.4, or correspondingly a Morse–Smale diffeomorphism of a non-orientable surface $N_{\mathrm{g}}$ given by Theorem 1.4, such that ${\mathcal{A}P}(f)={\mathcal{A}}$.
Then for every $n\in{\mathcal{A}P}_{odd}(f)$ and every $h\in C_{T}(S_{\mathrm{g}})$, respectively $h\in C_{T}(N_{\mathrm{g}})$, homotopic to $f$, we have $n\in{\rm Per}(h)$. For every $n\in{\mathcal{A}P}_{ev}(f)$ and every $h\in C_{T}(S_{\mathrm{g}})$, respectively $h\in C_{T}(N_{\mathrm{g}})$, homotopic to $f$, we have ($(\{n\}\cup\{\frac{n}{2}\})\cap{\rm Per}(h)\neq\emptyset$.
@@ 4. Final discussion and applications
In this section we describe deep relations of the considered concepts with Nielsen periodic point theory. In the next step, we give an estimate from below of the number of algebraically finite type mapping classes of surfaces homeomorphisms.
@@
We remind that the asymptotic Nielsen number of a self-map $f:M\to M$ of a compact manifold $M$, correspondingly the asymptotic generalized Lefschetz number of $f$, are defined as
@@
where the generalized Lefschetz number $L_{\Gamma}(f)$ is defined as an element of the group ring $\mathbb{Z}(\Gamma)$ of $\Gamma=\pi_{1}(T_{f})$ where $T_{f}$ is the mapping torus $f$ (cf. (1.4) [BoJia]).
For a surface without boundary, [BoJia, Thm 3.7] states the following: if $f:M\to M$ is a homeomorphism of a compact connected surface with $\chi(M)<0$ (orientable or not orientable), then
@@
where $\lambda$ is the largest stretching (expanding) factor of the pseudo-Anosov pieces in the Thurston canonical form of $f$. Lemma [BoJia, Lem. 3.6] adds that $\log\lambda={\bf h}(f)$ is equal to the topological entropy of the canonical representative of $f$.
@@
As a matter of fact in [BoJia, Thm 3.7] is shown more. Namely $N^{\infty}(f)=NI^{\infty}(f)$, where the latter is equal to lim $\limsup\sqrt[n]{NP_{n}(f)}$ and $NP_{n}(f)$ is so-called $n$-th Nielsen-Jiang prime periodic number (see [BoJia] for a definition, or [Je-Ma] for a longer exposition). This invariant has the property $NP_{n}(f)\leq|P_{n}(f)|$. It shows that in the case of a homeomorphisms of a surface $S_{\mathrm{g}}$ the condition $N^{\infty}(f)=\lambda>1$ implies that $P(f)={\underset{n=1}{\overset{\infty}{\cup}}}\,P^{n}(f)$, and also ${\rm Per}(f)$, are infinite. Since $N(f^{n})$ and $NP_{n}(f)$ are homotopy invariants, this yields that a Morse–Smale diffeomorphisms can only occur in the classes $T1$ or $T2$. In other words it provides one direction implication of Theorem 2.3: the set of the isotopy classes of Morse–Smale diffeomorphisms is contained in $T1\cup T2$.
@@ 4.1. Mapping classes
It is very useful to use the language of mapping classes in the study dynamics of homeomorphisms of surfaces. It is caused by the two factors. First, we are interested in homotopy properties of homeomorphisms of surfaces, second, the theory of mapping classes group is well-developed and several theorems we use are formulated in these terms.
@@
where ${\rm Homeo}^{+}(S_{\mathrm{g}})$ denotes the group of preserving orientation homeomorphisms of $S_{\mathrm{g}}$ and ${\rm Homeo}_{0}(S_{\mathrm{g}})$ its subgroup of homeomorphisms isotopic to the identity (cf. [Farb-Marg], [Par]). Thus the assignment $\phi\mapsto[\phi]\in\mathrm{M}od(S_{\mathrm{g}})$ is a surjection from ${\rm Homeo}^{+}(S_{\mathrm{g}})$ onto $\mathrm{M}od(S_{\mathrm{g}})$.
This allows us to define an action (a representation) of the group $\mathrm{M}od(S_{\mathrm{g}})$ in $H_{1}(S_{\mathrm{g}};\mathbb{Z})\subset H_{1}(S_{\mathrm{g}};\mathbb{R})$. More precisely, for a given class $[\phi]$ of $\phi\in{\rm Homeo}^{+}(S_{\mathrm{g}})$ we define $\Psi([\phi]):=H_{1}(\phi)\in Aut(H_{1}(S_{\mathrm{g}};\mathbb{Z})))$, where $H_{1}(\phi)$ is the induced homomorphism of $H_{1}(S_{\mathrm{g}};\mathbb{Z})$. It is well defined, i.e. does not depend on a choice of representative of the class, because the subgroup ${\rm Homeo}_{0}(S_{\mathrm{g}})$ acts trivially on $H_{1}(S_{\mathrm{g}};\mathbb{Z})$. In fact, $H_{1}(\phi)$ belongs to the group $Sp(2\mathrm{g},{\mathbb{Z}})$ of symplectic matrices of size $n$ according to the property (6).
Now we can formulate a classical fact (that we use below to study transversal maps) about the mapping class group (cf. [Farb-Marg, Proposition 7.3]).
@@
The homomorphism $\Psi:\mathrm{M}od(S_{\mathrm{g}})\to Sp(2\mathrm{g},\mathbb{Z})$ is surjective.
Another classical fact was shown by J. Nielsen by a direct geometrical consideration and a calculation of the characteristic polynomial (cf. [Nielsen]). Nowadays there are many ways of proving this statement.
@@
Combining the stated above facts we get the following.
@@
Every $f\in[h]$ of a homeomorphism $h:M\to M$ of orientable or non-orientable surface $M$ such that $N^{\infty}(h)=1$ is isotopic to a Morse–Smale diffeomorphism, thus it is quasi-unipotent and ${\mathcal{A}P}(f)$ is finite. Conversely, for every homeomorphism $f$ of $M$ which is homotopic to a Morse–Smale diffeomorphism we have $N^{\infty}(f)=1$.
@@
If $N^{\infty}(h)=1$ then by the Boju Jiang theorem there is not a pseudo-Anosov piece in the canonical Nielsen–Thurston form of $h$. Indeed, this theorem states that if $N^{\infty}(h)=\lambda$ then $\lambda$ is the largest stretching factor of the pseudo-Anosov pieces in the canonical form of $h$. However, for a pseudo-Anosov diffeomorphism we have $\lambda>1$ which leads to a contradiction if it would be such a piece.
Consequently, $[h]\in T1\cup T2$. Now, by the Nielsen theorem (Theorem 4.3) it is quasi-unipotent. By da Rocha’s Theorem 2.3 there is a Morse–Smale diffeomorphism $\tilde{h}\in[h]$.
Next, from Fact 2.5 it follows that ${\mathcal{A}P}(h)={\mathcal{A}P}(\tilde{h})$ is finite.
Finally, if $N^{\infty}(h)>1$ then $h$ and every $h^{\prime}\sim h$ has infinitely many periodic orbits which gives a contradiction if $h^{\prime}$ is Morse–Smale.
The last part, regarding entropy, follows from already quoted [BoJia, Lem. 3.6]. ∎
Let us remind that the mapping class group $\mathrm{M}od(S_{\mathrm{g}})$, can be also defined as the quotient ${\rm Diffeo}^{+}(S_{\mathrm{g}})/{\rm Diffeo}_{0}(S_{\mathrm{g}})$, (cf. [Farb-Marg]).
@@
For every symplectic or antisymplectic matrix $A\in Gl(2\mathrm{g},{\mathbb{Z}})$ there exists a transversal diffeomorphism $f\in DC_{T}(S_{\mathrm{g}})$, for which the induced automorphism $H_{1}(f)$ is equal to $A$.
@@
By Theorem 4.2 for every $A\in Sp(2\mathrm{g},\mathbb{Z})$ there exists a diffeomorphism $h$ of $S_{\mathrm{g}}$ for which the induced automorphism $H_{1}(h)$ is equal to $A$. To get a transversal diffeomorphism $f\in DC_{T}(S_{\mathrm{g}})$ homotopic to $h$, we use the fact that the set of all transversal self-diffeomorphisms of $S_{\mathrm{g}}$ is dense in ${\rm Diffeo}(S_{\mathrm{g}})$ by Theorem 3.6.
In the case of an antisymplectic matrix $A$ take any orientation-reversing homeomorphism $h\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$. Then the product $AH_{1}(h)$ is symplectic and by the first part of the proof induced by a transversal diffeomorphism $f^{\prime}$. Thus $f=f^{\prime}\circ h^{-1}$ induces $H_{1}(f)=A$ and again can be approximated by a transversal diffeomorphism. ∎
4.2. An estimate of the number of algebraically finite type classes
The formulas of Theorem 3.1 lead to an estimate from below of the number of algebraically finite type mapping classes for a fixed genus $\mathrm{g}$. Let us consider two homeomorphisms $f^{\prime},f^{\prime\prime}$ of a surface $M$. Observe that if they have different $k$-periodic expansions of Lefschetz numbers of iterations (cf. formula (2)) they could not be homotopic. As result the number of different expansions estimates the number of homotopy classes.
This observation and our construction allow us to find an estimate of the number of non-homotopic diffeomeorphisms in $T1\cup T2$.

### v2

@@ 3.1. Existence of periodic points for Morse–Smale diffeormophisms and transversal maps
To formulate consequences of our main result related to the existence of periodic points, we recall the notion of a transversal map.
Definition 3.5 ([5], [9]).
Let $f\colon{\mathcal{U}}\to M$ be a $C^{1}$-map of an open subset ${\mathcal{U}}$ of a manifold $M$. We say that $f$ is transversal if for any $m\in{{\mathbb{N}}}$ and $x\in P^{m}(f)$, $1\notin\sigma(Df^{m}(x))$, where $\sigma$ denotes the spectrum of the derivative $D$ of $f^{m}(x)$.
The set of all transversal maps ${\mathcal{U}}\to M$ is denoted by $C_{T}({\mathcal{U}},M)$ or $C_{T}(M)$ if ${\mathcal{U}}=M$. The main property of the class of transversal maps is given in the following theorem (see [5] if $M={\mathbb{R}}^{d}$, and [24] for the general case, also [14] for an exposition).
@@
The set $C^{\infty}_{T}({\mathcal{U}},M)$ is generic in $C^{0}({{\mathcal{U}}},M)$, i.e. it is the intersection $C^{\infty}_{T}({\mathcal{U}},M)=\bigcap_{n=1}^{\infty}G_{n}$ where $G_{n}$ is open and dense in $C^{0}({{\mathcal{U}}},M)$ In particular, every map $f\colon{\mathcal{U}}\to M$ is homotopic to a transversal map $h\colon{\mathcal{U}}\to M$.
A geometric property of a transversal map is as follows.
@@
Let $M$ be a closed manifold. For any $f\in C_{T}(M)$ and every $m\in{\mathbb{N}}$, the set $P^{m}(f)$ consists of isolated points.
It follows easily from the definition that Morse–Smale diffeomorphisms belong to $C_{T}(M)$ (cf. [27], [9]). We denote the set of Morse–Smale diffeomorphisms by $DC_{MS}(M)$.
Finally, all mappings in $C_{T}(M)$, thus in $DC_{MS}(M)$, have the following geometric property, which states that non-vanishing of $a_{n}(f)$, $n$-odd, implies that $P_{n}(f)\neq\emptyset$, i.e. $n\in{\rm Per}(f)$.
Proposition 3.8 (see [14, Corollary 3.3.10] for an exposition, cf. [5, 9] for the case of Morse–Smale diffeomorphisms).
@@
Let $f\colon M\to M$ be a transversal map.
If $a_{n}(f)\neq 0\,$ then $\;\begin{cases}P_{n}(f)\cup P_{\frac{n}{2}}(f)\neq\emptyset\;\,{\text{if}}\;n\;{\text{is even}},\cr P_{n}(f)\neq\emptyset\;\;{\text{if}}\;n\;{\text{is odd}}.\end{cases}$
The above proposition allows us to formulate the analytical (dynamical) consequence of Theorem 1.3. Let ${\mathcal{A}P}_{odd}(f)={\mathcal{A}P}(f)\cap(2{\mathbb{N}}-1)$, and ${\mathcal{A}P}_{even}(f)={\mathcal{A}P}(f)\cap(2{\mathbb{N}})$.
@@
Let ${\mathcal{A}}\subset{\mathbb{N}}$ be a finite subset of natural numbers, and $f\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ be a preserving or reversing orientation Morse–Smale diffeomorphism of orientable surface of genus $\mathrm{g}$, or correspondingly a Morse–Smale diffeomorphism of a non-orientable surface $N_{\mathrm{g}}$ given by Theorem 1.3 such that ${\mathcal{A}P}(f)={\mathcal{A}}$.
Moreover, let $h\in C_{T}(S_{\mathrm{g}})$, respectively $h\in C_{T}(N_{\mathrm{g}})$, be homotopic to $f$ (in particular, $h$ can be equal to $f$). Then, $n\in{\rm Per}(h)$ for every $n\in{\mathcal{A}P}_{odd}(f)$ and ($(\{n\}\cup\{\frac{n}{2}\})\cap{\rm Per}(h)\neq\emptyset$ for every $n\in{\mathcal{A}P}_{even}(f)$.
@@ 4. Final discussion and applications
In this section, we describe deep relations of the considered concepts with Nielsen periodic point theory. In the next step, we give an estimate from below of the number of algebraically finite type mapping classes of surface homeomorphisms.
@@
We recall that the asymptotic Nielsen number of a self-map $f\colon M\to M$ of a compact manifold $M$, correspondingly the asymptotic generalized Lefschetz number of $f$, are defined as
@@
where the generalized Lefschetz number $L_{\Gamma}(f)$ is defined as an element of the group ring ${\mathbb{Z}}(\Gamma)$ of $\Gamma=\pi_{1}(T_{f})$ where $T_{f}$ is the mapping torus $f$ (cf. (1.4) [15]).
For a surface without boundary, [15, Thm 3.7] states the following: If $f\colon M\to M$ is a homeomorphism of a compact connected surface with $\chi(M)<0$ (orientable or not orientable), then
@@
where $\lambda$ is the largest stretching (expanding) factor of the pseudo-Anosov pieces in the Thurston canonical form of $f$. Lemma [15, Lem. 3.6] adds that $\log\lambda={\bf h}(f)$ is equal to the topological entropy of the canonical representative of $f$.
@@
In fact, a stronger result is established in [15, Thm 3.7]. Namely, $N^{\infty}(f)=NI^{\infty}(f)$, where the latter is equal to lim $\limsup\sqrt[n]{NP_{n}(f)}$, and $NP_{n}(f)$ is the so-called $n$-th Nielsen-Jiang prime periodic number (see [15] for a definition, or [14] for a longer exposition). This invariant has the property $NP_{n}(f)\leq|P_{n}(f)|$. It shows that in the case of a homeomorphism of a surface $S_{\mathrm{g}}$, the condition $N^{\infty}(f)=\lambda>1$ implies that $P(f)={\underset{n=1}{\overset{\infty}{\bigcup}}}\,P^{n}(f)$, and also ${\rm Per}(f)$, are infinite. Since $N(f^{n})$ and $NP_{n}(f)$ are homotopy invariants, this yields that a Morse–Smale diffeomorphisms can only occur in the classes $T1$ or $T2$. In other words, it establishes an implication in one direction of Theorem 2.3: The set of the isotopy classes of Morse–Smale diffeomorphisms is contained in $T1\cup T2$.
@@ 4.1. Mapping classes
It is very useful to use the language of mapping classes in the study dynamics of homeomorphisms of surfaces. It is caused by two factors. First, we are interested in homotopy properties of homeomorphisms of surfaces. Second, the theory of mapping classes groups is well-developed, and several theorems we use are formulated in these terms.
@@
where ${\rm Homeo}^{+}(S_{\mathrm{g}})$ denotes the group of preserving orientation homeomorphisms of $S_{\mathrm{g}}$, and ${\rm Homeo}_{0}(S_{\mathrm{g}})$ its subgroup of homeomorphisms isotopic to the identity (cf. [7], [22]). Thus, the assignment $\phi\mapsto[\phi]\in\mathrm{Mod}(S_{\mathrm{g}})$ is a surjection from ${\rm Homeo}^{+}(S_{\mathrm{g}})$ onto $\mathrm{Mod}(S_{\mathrm{g}})$.
This allows us to define an action (a representation) of the group $\mathrm{Mod}(S_{\mathrm{g}})$ in $H_{1}(S_{\mathrm{g}};{\mathbb{Z}})\subset H_{1}(S_{\mathrm{g}};{\mathbb{R}})$. More precisely, for a given class $[\phi]$ of $\phi\in{\rm Homeo}^{+}(S_{\mathrm{g}})$, we define $\Psi([\phi]):=H_{1}(\phi)\in Aut(H_{1}(S_{\mathrm{g}};{\mathbb{Z}})))$, where $H_{1}(\phi)$ is the induced homomorphism of $H_{1}(S_{\mathrm{g}};{\mathbb{Z}})$. It is well defined, i.e. it does not depend on a choice of representative of the class, because the subgroup ${\rm Homeo}_{0}(S_{\mathrm{g}})$ acts trivially on $H_{1}(S_{\mathrm{g}};{\mathbb{Z}})$. In fact, $H_{1}(\phi)$ belongs to the group $Sp(2\mathrm{g},{\mathbb{Z}})$ of symplectic matrices of size $n$ according to property (6).
Now we can formulate a classical fact (that we use below to study transversal maps) about the mapping class group (cf. [7, Proposition 7.3]).
@@
The homomorphism $\Psi\colon\mathrm{Mod}(S_{\mathrm{g}})\to Sp(2\mathrm{g},\mathbb{Z})$ is surjective.
Another classical fact was shown by J. Nielsen by a direct geometrical consideration and a calculation of the characteristic polynomial (cf. [21]). Nowadays there are many ways of proving this statement.
@@
Combining the facts stated above, we get the following.
@@
Every $f\in[h]$ of a homeomorphism $h\colon M\to M$ of orientable or non-orientable surface $M$ such that $N^{\infty}(h)=1$ is isotopic to a Morse–Smale diffeomorphism, thus it is quasi-unipotent and ${\mathcal{A}P}(f)$ is finite. Conversely, for every homeomorphism $f$ of $M$ which is homotopic to a Morse–Smale diffeomorphism, we have $N^{\infty}(f)=1$.
@@
If $N^{\infty}(h)=1$, then by the Boju Jiang theorem there is not a pseudo-Anosov piece in the canonical Nielsen–Thurston form of $h$. Indeed, this theorem states that if $N^{\infty}(h)=\lambda$, then $\lambda$ is the largest stretching factor of the pseudo-Anosov pieces in the canonical form of $h$. However, for a pseudo-Anosov diffeomorphism we have $\lambda>1$, which leads to a contradiction if it would be such a piece.
Consequently, $[h]\in T1\cup T2$. Now, by the Nielsen theorem (Theorem 4.3), it is quasi-unipotent. By da Rocha’s Theorem 2.3, there is a Morse–Smale diffeomorphism $\tilde{h}\in[h]$.
Next, from Fact 2.5, it follows that ${\mathcal{A}P}(h)={\mathcal{A}P}(\tilde{h})$ is finite.
Finally, if $N^{\infty}(h)>1$, then $h$ and every $h^{\prime}\sim h$ has infinitely many periodic orbits, which gives a contradiction if $h^{\prime}$ is a Morse–Smale diffeomorphism.
The last part regarding entropy follows from the already quoted [15, Lem. 3.6]. ∎
Let us remind that the mapping class group $\mathrm{Mod}(S_{\mathrm{g}})$ can be also defined as the quotient ${\rm Diffeo}^{+}(S_{\mathrm{g}})/{\rm Diffeo}_{0}(S_{\mathrm{g}})$, (cf. [7]).
@@
For every symplectic or antisymplectic matrix $A\in Gl(2\mathrm{g},{\mathbb{Z}})$, there exists a transversal map $f\in C_{T}(S_{\mathrm{g}})$ for which the induced automorphism $H_{1}(f)$ is equal to $A$.
@@
By Theorem 4.2, for every $A\in Sp(2\mathrm{g},{\mathbb{Z}})$, there exists a diffeomorphism $h$ of $S_{\mathrm{g}}$ for which the induced automorphism $H_{1}(h)$ is equal to $A$. Next, we replace $h$ by a transversal map $f\in C_{T}(S_{\mathrm{g}})$ homotopic to $h$ using Theorem 3.6.
In the case of an antisymplectic matrix $A$, take any orientation-reversing homeomorphism $h\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$. Then, the product $AH_{1}(h)$ is symplectic, and by the first part of the proof induced by a transversal map $f^{\prime}$. Thus, $f=f^{\prime}\circ h^{-1}$ induces $H_{1}(f)=A$, and again can be approximated by a transversal map. ∎
4.2. An estimate of the number of conjugacy classes of algebraically finite type mapping classes
The formulas of Theorem 3.1 lead to an estimate from below of the number of conjugacy classes of algebraically finite type, mapping classes for a fixed genus $\mathrm{g}$. Let us consider two homeomorphisms $f^{\prime},f^{\prime\prime}$ of a surface $M$. If they are conjugated, then they induce conjugated homomorphisms on each homology group, in particular $\Psi(f^{\prime})=H_{1}(f^{\prime})$ and $\Psi(f^{\prime\prime})=H_{1}(f^{\prime\prime})$ are conjugated matrices in $Sp(2\mathrm{g},\mathbf{Z})$. Since conjugated matrices have the same traces, $f^{\prime}$ and $f^{\prime\prime}$ have the same periodic expansions of sequences of Lefschetz numbers of iterations. As a result, the number of different periodic expansions estimates the number of conjugacy classes of given elements in $\mathrm{Mod}(M)$.
This observation and our construction allow us to find an estimate of the number of conjugacy classes of homotopy classes of diffeomorphisms in $T1\cup T2$.